OpenStudy (blu_seaa):

Which number in scientific notation is one thousand times larger than 3.7 x 103 ? A) 3.7 x 104 B) 3.7 x 105 C) 3.7 x 106 D) 3.7 x 107

OpenStudy (mariakershaw):

ok wat you do is mutlpliy each one then compare and see which is bigger/ greater @Blu_Seaa

OpenStudy (mariakershaw):

including 3.7*103 Kk.

OpenStudy (mariakershaw):

brb @Blu_Seaa

OpenStudy (blu_seaa):

Kk

OpenStudy (mariakershaw):

ok so did you multiply all of them and compare which should give you ur answer

OpenStudy (mariakershaw):

so wat is 3.7*104=

OpenStudy (blu_seaa):

384.8

OpenStudy (mariakershaw):

ok then wat is 3.7*105=

OpenStudy (blu_seaa):

388.5

OpenStudy (mariakershaw):

then wat is 3.7*106=

OpenStudy (blu_seaa):

392.2

OpenStudy (mariakershaw):

then wat is 3.7*107 srry its alot xD

OpenStudy (blu_seaa):

395.9

OpenStudy (mariakershaw):

ok finally wat is 3.7*103 equal

OpenStudy (blu_seaa):

381.1

OpenStudy (mariakershaw):

ok so out 395.9,392.2,388.5,and 384.8 which is bigger that 381.1

OpenStudy (blu_seaa):

395.9

OpenStudy (mariakershaw):

so which equation was that and that is yur answer =)

OpenStudy (blu_seaa):

D, thanks again. I just wanted to double check and make sure I got it.

OpenStudy (mariakershaw):

ok

OpenStudy (mariakershaw):

then when yur ready post yur next one